# Do global forecasting models require frequent retraining?

### v1

@@ 3.2.2 Deep learning models
Multi-Layer Perceptron (MLP) is a feedforward neural network consisting of an input layer, one or more hidden layers, and an output layer. Each layer applies a non-linear activation function to capture complex relationships in the data (Rosenblatt, 1958). MLPs are a versatile and very efficient solution. For this reason, many deep learning models specifically created for time series forecasting are MLP-based.
Recurrent Neural Networks (RNN) are designed to model sequential data by maintaining a hidden state that captures temporal dependencies. Standard RNNs, however, suffer from vanishing gradient problems, limiting their ability to learn long-term dependencies (Cho et al., 2014). Variants such as Long Short-Term Memory (LSTM) networks and Gated Recurrent Units (GRUs) address this limitation and are widely used for time series forecasting (Hochreiter & Schmidhuber, 1997) 666 In our work, we used only the LSTM version but, for the cost of computation, it was trained only on the VN1 weekly dataset. .
@@
Neural Hierarchical Interpolation for Time Series (NHITS) builds upon the success of NBEATS by incorporating hierarchical interpolation mechanisms to better capture time-hierarchies in time series data. Multi-rate input pooling, hierarchical interpolation, and backcast residual connections together induce the specialization of the additive predictions in different signal bands, reducing memory and computational time, thus improving the architecture’s parsimony and accuracy (Challu et al., 2022).
Deep learning models typically do not require the extensive feature engineering step needed by their machine learning counterpart, since they create such features (like lags and rolling averages) internally. Nevertheless, they usually are more difficult to train since they have much more hyperparameters to select affecting the forecasting performance (Smyl, 2020). We trained the global deep learning models by adding static, calendar, and external features only, and relying on top competitions’ solutions to set the hyperparameters’ values.
@@ 3.3 Retrain scenarios
To answer our research question we explored several possible retraining scenarios. A retrain scenario or retrain window, $r$, is identified as a positive integer representing the frequency at which the model is re-trained or updated. Essentially, it represents how many data points need to be passed before a new training step is performed. The retrain scenarios strongly depend on the frequency of the dataset considered because the frequency drives both the forecast horizon and the business review periods. Therefore, we defined different set of retrain scenarios based on the frequencies of our datasets:
@@
Note that we are training global models, hence the training data set is composed by the training set of each time series in the dataset. Since the datasets we considered are all aligned, this implies that, at every retrain period, we fit the model on a new training data set containing $r$ new observations for each time series (where $r$ is the chosen retrain scenario). Moreover, note also that we are considering only two different forms of model update: a model can either be completely updated when retraining is performed, or used as is to produce forecasts for $r$ subsequent periods. Therefore, we are not examining the effect of hyperparameter tuning within each retrain scenario, given the expensive computational cost of this process and minor changes expected.
The scenario $r=1$ is the so-called continuous retraining, and it is the most expensive since it implies that the model is retrained every new observation. Theoretically, but usually also practically, it should be the most accurate forecasting scenario, because the model used to predict has been trained on all the available data points up to that moment. For this reason, we considered this scenario as the benchmark both in terms of forecasting accuracy and computational cost. For daily data only, however, the benchmark scenario is $r=7$, because it is not common to retrain a global model every day, and usually the update is performed once a week. The scenario $r=T$ is the no retraining scenario, meaning that the forecasting model is fitted just once on the initial training set, then it is used to produce the forecasts for the entire test set $T$. It is the lowest computationally expensive but it should also be the lowest accurate. All the other scenarios such that $1<r<T$ are considered as periodic updating strategies. Both the accuracy and the computing time should be non-increasing functions of $r$.
@@ 3.4 Evaluation setup
In time series forecasting, out-of-sample testing is crucial for assessing how well a model can generalize to unseen data. This is particularly important because patterns in historical data may not fully represent future trends, and unexpected changes in the data may occur (Tashman, 2000).
As suggested by Bergmeir & Benítez (2012), rolling origin evaluation is the most widely used and correct method for conducting such tests, offering a systematic way to evaluate forecasting models over multiple iterations.
@@
The rolling origin evaluation process begins by dividing the time series into two parts: a training set, which is used to build the model, and a test set, which is used for evaluation. The test set always follows the training set chronologically to maintain the temporal order of the data. At each step, the model is trained on the training set and generates forecasts for a specified number of future points, the forecast horizon $h$. These forecasts are then compared to the actual values in the test set using a chosen evaluation metric (see Section 3.5). What makes rolling origin evaluation unique is its iterative nature. After each forecasting step, the forecast origin (the last point in the training set) is shifted forward by a fixed number of periods, the step size. The model is then retrained using the updated training set, which may either expand to include all available data up to the new forecast origin or remain fixed to a specific window size. This process is repeated until the test set is fully utilized, and the overall performance of the model is calculated by averaging the results across all the iterations.
@@
Compared to the fixed origin evaluation setup, that allows for a single evaluation step, the rolling origin evaluation is preferable for its ability to provide a more comprehensive assessment of a model’s performance under different conditions, such as varying seasonal patterns, level shifts, or data trends. By simulating multiple forecast cycles, indeed, this method reduces the risk of bias that can arise from evaluating a model a single time and at a single forecast origin Bergmeir & Benítez (2012). This approach is beneficial in industries like retail and supply chain management, where forecasts need to be continuously updated to reflect real-time changes. By simulating multiple scenarios and evaluating a model’s performance across different cycles, rolling origin evaluation provides valuable insights into the robustness and reliability of the forecasting methods adopted.
@@
In practice, the rolling origin evaluation setup can be adjusted based on the data and forecasting goals. For instance, the training set can be fixed to a rolling window of the most $n$ recent observations (fixed window), or it can expand to include the full historical data (expanding window) Bergmeir & Benítez (2012). In real-life applications, most forecast practitioners use this second option, especially when the length of the time series is small (Petropoulos & et al., 2022). In our study, we adopted the expanding window approach, because we are interested in simulating a forecasting experiment as close as possible to the business reality. Moreover, it is also the only reasonable option when dealing with short time series, like the case for the weekly VN1 datasets.
@@
Table 2: The retraining scenarios associated with each dataset. The test windows and the forecasting horizons are chosen based on the frequency of the dataset.
Dataset Frequency Retraining Scenarios (r) Test Window (T) Horizon (h)
M5 Daily (7) 7, 14, 21, 30, 60, 90, 120, 150, 180, 364 364 28
VN1 Weekly (52) 1, 2, 3, 4, 6, 8, 10, 13, 26, 52 52 13
@@
The size of the test set, the length of the forecast horizon, and the step size for shifting the forecast origin are also customizable parameters that depend on the specific use case. Table 2 shows the parameters selected in our experiments. We set the size of the test set to cover at least one complete year so that the evaluation of the different scenarios does not depend on intra-year variations, such as any specific season or period of the year. We selected the forecast horizons based on the type of business decisions that usually daily and weekly forecasts support, i.e. operational vs strategic planning. Finally, the step size was set always equal to one to maximize the number of evaluations of each scenario.
@@ 3.5 Performance metrics
The accuracy evaluation of point predictive models is a controversial topic in time series forecasting: many metrics are available to capture models’ performances but no consensus has been reached in the literature on whether one metric is better than others (Hewamalage et al., 2023). However, since we are dealing with demand forecasting at SKU-level, implying that data intermittency is very much likely, all the metrics based on absolute or percentage errors are not optimal since they optimize for the median (Kolassa, 2020). Moreover, being the scale possibly different among the series, a scaled accuracy measure is preferable. For this reason, we considered the Root Mean Squared Scaled Error (RMSSE), proposed by Hyndman & Koehler (2006), to evaluate the point forecast accuracy of the models. The RMSSE is defined as:
@@
The RMSSE measures the relative prediction accuracy of a forecasting method by comparing the mean squared errors of the prediction and the observed value against the mean squared errors of the seasonal naive model. This metric was the official metric used to assess the performance of models in the M5 competition (with $s=1$) Makridakis et al. (2022b), which is one of the datasets we used in our experiments. Lower RMSSE values suggest higher model accuracy.
@@
Our work is not limited to the evaluation of point forecasting accuracy, but we aim also to assess the probabilistic performance of the models throughout each different retrain scenario. However, since the machine learning and deep learning models adopted for the experiments do not output density (or probabilistic) predictions by default, we relied on Conformal Inference as a general framework to build prediction intervals. Conformal Inference is a powerful tool that allows quantification of the uncertainty associated with out-of-sample predictions simply using a validation set on a point forecaster model to produce intervals at the desired levels (Vovk et al., 2005). It was originally proposed for non-temporal data, because the data exchangeability assumption, nevertheless in recent years it has been extended to be used also in the context of time series data (Stankeviciute et al., 2021). In particular, given its guaranteed coverage, distribution-free, model-agnostic, computational efficiency, and low dataset size requirements, Conformal Inference is likely the only tool available to evaluate the uncertainty in a global forecasting setting where different models are compared among several datasets and scenarios.
@@
Once the prediction intervals are obtained, we used the Multi-Quantile Loss to comprehensively measure the goodness of the probabilistic predictions. The Quantile Loss (also known as Pinball Loss) and the Multi-Quantile Loss are defined as:
$\text{QL}=\frac{1}{h}\sum_{t=n+1}^{n+h}\left(q\cdot(y_{t}-\hat{y}_{t})\cdot\mathbb{I}_{y_{t}\geq\hat{y}_{t}}+(1-q)\cdot(\hat{y}_{t}-y_{t})\cdot\mathbb{I}_{y_{t}<\hat{y}_{t}}\right)\mbox{,}$ (4)
$\text{MQL}=\frac{1}{\mathcal{Q}}\sum_{q\in\mathcal{Q}}\text{QL}(q)\mbox{.}$ (5)
The QL, being a proper scoring rule, allows for the correct evaluation of the probabilistic forecasts (Kolassa, 2016). Moreover, the (Weighted) MQL was the official metric used to evaluate the overall performance of the forecasting methods during the M5 Uncertainty competition (Makridakis et al., 2022c).
@@
Note that we considered the median and 6 central prediction intervals, namely 60%, 70%, 80%, 90%, 95% and 99%, for a total of 13 different quantiles. The median and the 60% and 70% central prediction intervals provide a good description of the center of the forecast distribution, while the 90%, 95%, and 99% give information about its tails, that in retail demand forecasting problems are fundamental to determine the appropriate safety stock levels (Barrow & Kourentzes, 2016). These quantiles provide sufficient information about the uncertainty of the forecasts and allow for the effective description of the whole distribution.
@@
Note also that we computed conformal prediction intervals on a validation set at least twice as long as the forecast horizon. This was the main reason that forced us to reduce the number of time series in the datasets, that is to be able to obtain a reliable estimate of the model uncertainty through prediction intervals.
@@
Finally, we addressed the problem of evaluating the cost associated to producing the forecasts by calculating the Computing Time (CT). CT is defined as the time in seconds required for training and predicting the next $h$ time steps ahead with the model. Among the possible measures of complexity, such as the number of parameters, number of iterations, model’s depth, etc, CT is the most directly related to the monetary cost of forecasting since the forecasts are usually produced through some sort of cloud computing service that relies on pay-as-you-go fares (Spiliotis & Petropoulos, 2024). Hence, CT is analyzed for each forecasting model and retrain scenario combination. Lower CT values imply lower forecasting costs.
@@
Note that we related each evaluation metric result to the baseline retraining scenario, depending on the frequency of the dataset, to be able to easily compare the performance among different models and retrain windows. Furthermore, to statistically verify our research question we tested the scenarios’ results using the Friedman-Nemenyi test for multiple comparisons (Demšar, 2006).

### v2

@@ 3.2.2 Deep learning models
Multi-Layer Perceptron (MLP) is a feedforward neural network consisting of an input layer, one or more hidden layers, and an output layer. Each layer applies a non-linear activation function to capture complex relationships in the data (Rosenblatt, 1958). MLPs are a versatile and very efficient solution. For this reason, many deep learning models specifically created for time series forecasting are MLP-based. MLP is also considered a solid benchmark for deep learning global forecasting models (Godahewa et al., 2021b).
Recurrent Neural Networks (RNN) are designed to model sequential data by maintaining a hidden state that captures temporal dependencies. Standard RNNs, however, suffer from vanishing gradient problems, limiting their ability to learn long-term dependencies (Cho et al., 2014). Variants such as Long Short-Term Memory (LSTM) networks and Gated Recurrent Units (GRUs) address this limitation and are widely used for time series forecasting (Hochreiter & Schmidhuber, 1997) 666 In our work, we used only the LSTM version, but, for the cost of computation, it was trained only on the VN1 weekly dataset. .
@@
Neural Hierarchical Interpolation for Time Series (NHITS) builds upon the success of NBEATS by incorporating hierarchical interpolation mechanisms to better capture time hierarchies in time series data. Multi-rate input pooling, hierarchical interpolation, and backcast residual connections together induce the specialization of the additive predictions in different signal bands, reducing memory and computational time, thus improving the architecture’s parsimony and accuracy (Challu et al., 2022).
Deep learning models typically do not require extensive feature engineering step needed by their machine learning counterpart, since they create such features (like lags and rolling averages) internally. Nevertheless, they usually are more difficult to train since they have many more hyperparameters to select affecting the forecasting performance (Smyl, 2020). We trained the global deep learning models by adding static, calendar, and external features only, and relying on the top competitions’ solutions to set the hyperparameters’ values.
@@ 3.3 Retrain scenarios
To answer our research question, we explored several possible retraining scenarios. A retrain scenario or retrain window, $r$, is identified as a positive integer representing the frequency at which the model is re-trained or updated. Essentially, it represents how many data points need to be passed before a new training step is performed. The retrain scenarios strongly depend on the frequency of the dataset considered because the frequency drives both the forecast horizon and the business review periods. Therefore, we defined different set of retrain scenarios based on the frequencies of our datasets:
@@
Note that we are training global models, hence the training dataset is composed by the training set of each time series in the dataset. Since the datasets we considered are all aligned, which implies that, at every retrain period, we fit the model on a new training data set containing $r$ new observations for each time series (where $r$ is the chosen retrain scenario). Moreover, note also that we are considering only two different forms of model update: a model can either be completely updated when retraining is performed, or used as is to produce forecasts for $r$ subsequent periods. Therefore, we are not examining the effect of hyperparameter tuning within each retrain scenario, given the expensive computational cost of this process and minor changes expected.
The scenario $r=1$ is the so-called continuous retraining, and it is the most expensive since it implies that the model is retrained for every new observation. Theoretically, but usually also practically, it should be the most accurate forecasting scenario, because the model used to predict has been trained on all the available data points up to that moment. For this reason, we considered this scenario as the benchmark both in terms of forecasting accuracy and computational cost. For daily data only, however, the benchmark scenario is $r=7$, because it is not common to retrain a global model every day, and usually the update is performed once a week. The scenario $r=T$ is the no retraining scenario, meaning that the forecasting model is fitted just once on the initial training set, then it is used to produce the forecasts for the entire test set $T$. It is the lowest computationally expensive but it should also be the lowest accurate. All the other scenarios such that $1<r<T$ are considered as periodic updating strategies. Both the accuracy and the computing time should be non-increasing functions of $r$.
@@ 3.4 Evaluation setup
In time series forecasting, out-of-sample testing is crucial for assessing how well a model can generalize to unseen data. This is particularly important because patterns in historical data may not fully represent future trends, and unexpected changes in the data may occur (Tashman, 2000). As suggested by Bergmeir & Benítez (2012), rolling origin evaluation is the most widely used and correct method for conducting such tests, offering a systematic way to evaluate forecasting models over multiple iterations. The rolling origin evaluation process begins by dividing the time series into two parts: a training set, which is used to build the model, and a test set, which is used for evaluation. The test set always follows the training set chronologically to maintain the temporal order of the data. At each step, the model is trained on the training set and generates forecasts for a specified number of future points, the forecast horizon $h$. These forecasts are then compared to the actual values in the test set using a chosen evaluation metric (see Section 3.5). What makes rolling origin evaluation unique is its iterative nature. After each forecasting step, the forecast origin (the last point in the training set) is shifted forward by a fixed number of periods, the step size. The model is then retrained using the updated training set, which may either expand to include all available data up to the new forecast origin or remain fixed to a specific window size. This process is repeated until the test set is fully utilized, and the overall performance of the model is calculated by averaging the results across all the iterations.
@@
Compared to the fixed origin evaluation setup, which allows for a single evaluation step, the rolling origin evaluation is preferable for its ability to provide a more comprehensive assessment of a model’s performance under different conditions, such as varying seasonal patterns, level shifts, or data trends. By simulating multiple forecast cycles, indeed, this method reduces the risk of bias that can arise from evaluating a model a single time and at a single forecast origin (Bergmeir & Benítez, 2012). This approach is beneficial in industries like retail and supply chain management, where forecasts need to be continuously updated to reflect real-time changes. By simulating multiple scenarios and evaluating a model’s performance across different cycles, rolling origin evaluation provides valuable insights into the robustness and reliability of the forecasting methods adopted.
@@
In practice, the rolling origin evaluation setup can be adjusted based on the data and forecasting goals. For instance, the training set can be fixed to a rolling window of the most $n$ recent observations (fixed window), or it can expand to include the full historical data (expanding window) (Bergmeir & Benítez, 2012). In real-life applications, most forecast practitioners use this second option, especially when the length of the time series is small (Petropoulos & et al., 2022). In our study, we adopted the expanding window approach because we are interested in simulating a forecasting experiment as close as possible to the business reality. Moreover, it is also the only reasonable option when dealing with short time series, like the case for the weekly VN1 datasets.
@@
Table 2: Retraining scenarios, test windows, and forecast horizons for each dataset. The setup is based on the respective time series frequency.
M5 VN1
Training Window ($w$) $\geq 730$ $\geq 157$
Test Window ($T$) 364 52
Horizon ($h$) 28 13
Step size ($p$) 1 1
Retraining Scenarios ($r$) {7, 14, 21, 30, 60, 90, 120, 150, 180, 364} {1, 2, 3, 4, 6, 8, 10, 13, 26, 52}
@@
The size of the test set $T$, the length of the forecast horizon $h$, and the step size $p$, for shifting the forecast origin, are also customizable parameters that depend on the specific use case. Table 2 shows the parameters selected in our experiments. Following Spiliotis & Petropoulos (2024), we set the size of the test set to cover at least one complete year so that the evaluation of the different scenarios does not depend on intra-year variations, such as any specific season or period of the year. We selected the forecast horizons based on the type of business decisions that usually daily and weekly forecasts support, i.e. operational vs strategic planning. Finally, the step size was always set equal to one to maximize the number of evaluations of each scenario.
@@ 3.5 Performance metrics
The accuracy evaluation of point predictive models is a controversial topic in time series forecasting: many metrics are available to capture models’ performances, but no consensus has been reached in the literature on whether one metric is better than others (Hewamalage et al., 2023). However, since we are dealing with demand forecasting at SKU-level, implying that data intermittency is very much likely, all the metrics based on absolute or percentage errors are not optimal since they optimize for the median (Kolassa, 2020). Moreover, being the scale possibly different among the series, a scaled accuracy measure is preferable. For this reason, we considered the Root Mean Squared Scaled Error (RMSSE), proposed by Hyndman & Koehler (2006), to evaluate the point forecast accuracy of the models. The RMSSE is defined as:
@@
The RMSSE measures the relative prediction accuracy of a forecasting method by comparing the mean squared errors of the prediction and the observed value against the mean squared errors of the seasonal naive model. This metric was the official metric used to assess the performance of models in the M5 competition (with $s=1$, hence using the naive model as a benchmark) (Makridakis et al., 2022b). Lower RMSSE values suggest higher model accuracy.
@@
Our work is not limited to the evaluation of point forecasting accuracy, but we also aim to assess the probabilistic performance of the models throughout each different retrain scenario. However, since the machine learning and deep learning models adopted for the experiments do not output density (or probabilistic) predictions by default, we relied on Conformal Inference as a general framework to build prediction intervals. Conformal Inference is a powerful tool that allows quantification of the uncertainty associated with out-of-sample predictions simply using a validation set on a point forecaster model to produce intervals at the desired levels (Vovk et al., 2005). It was originally proposed for non-temporal data, because the data exchangeability assumption, nevertheless, in recent years it has been extended to be used also in the context of time series data (Stankeviciute et al., 2021). In particular, given its guaranteed coverage, distribution-free, model-agnostic, computational efficiency, and low dataset size requirements, Conformal Inference is likely the only tool available to evaluate the uncertainty in a global forecasting setting where different models are compared among several datasets and scenarios. One may prefer, instead, to train models using quantile losses to directly yield quantile forecasts, but this would require training a single model for each quantile. Even if it may be a valuable way of modeling uncertainty, as demonstrated by the M5 Uncertainty competition (Makridakis et al., 2022c; Lainder & Wolfinger, 2022), it has the huge disadvantage of being very computationally expensive, hence negatively impacting on the sustainability of forecasting applications in large-scale environments. Moreover, even if multi-quantile losses can provide multiple quantiles, and prediction intervals, they usually are uncalibrated, meaning that the actual frequency of observations falling within the interval does not align with the confidence level associated with it. Therefore, for these multiple reasons, we decided to adopt Conformal Inference as general tool to estimate quantile forecasts.
@@
Once the quantiles are obtained, we used a scaled version of the Multi-Quantile Loss to comprehensively measure the goodness of the probabilistic predictions. The scaled Quantile Loss, also known as (scaled) Pinball Loss, and the scaled Multi-Quantile Loss are defined as:
$\text{SQL}=\frac{1}{h}\frac{\sum_{t=n+1}^{n+h}\left(q\cdot(y_{t}-\hat{y}_{t})\cdot\mathbb{I}_{y_{t}\geq\hat{y}_{t}}+(1-q)\cdot(\hat{y}_{t}-y_{t})\cdot\mathbb{I}_{y_{t}<\hat{y}_{t}}\right)}{\frac{1}{n-s}\sum_{t=s+1}^{n}|y_{t}-y_{t-s}|}\mbox{,}$ (4)
$\text{SMQL}=\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\text{SQL}(q)\mbox{,}$ (5)
where $\mathcal{Q}$ is the set of possible quantiles. The QL, being a proper scoring rule, allows for the correct evaluation of the probabilistic forecasts (Kolassa, 2016). The scaling factor is the in-sample, one-step-ahead mean absolute error of the seasonal Naive model. The scaled MQL (weighted and with $s=1$) was also the official metric used to evaluate the overall performance of the forecasting methods during the M5 Uncertainty competition (Makridakis et al., 2022c). For daily data we used $s=7$, because $s=1$ is arguably an irrelevant benchmark in daily retail data, which typically exhibits strong day-of-week seasonality (Fildes et al., 2022). Instead, we set $s=1$ for weekly data.
@@
Note that we considered a total of 14 different quantiles, $q$, of probability levels {0.005, 0.025, 0.050, 0.100, 0.150, 0.200, 0.025, 0.750, 0.800, 0.850, 0.900, 0.950, 0.975, 0.995}. Therefore, we studied the 50% 60%, 70%, 80%, 90%, 95%, and 99% central prediction intervals. The 50% and 60% central prediction intervals provide a good description of the center of the forecast distribution, while the 90%, 95%, and 99% give more information about its tails, which in retail demand forecasting problems are fundamental to determine the appropriate safety stock levels (Barrow & Kourentzes, 2016). These quantiles provide sufficient information about the uncertainty of the forecasts and allow for the effective description of the whole distribution.
@@
Note also that we computed conformal prediction intervals on a validation set at least twice as long as the forecast horizon. This was the main reason that forced us to reduce the number of time series in the datasets, that is to be able to obtain a reliable estimate of the model uncertainty through prediction intervals 777 We used a validation set four times larger than the forecast horizon for the M5 dataset and twice the forecast horizon for the VN1 dataset. This limited the number of time series from 30,490 to 28,298 for the M5 dataset. .
@@
Finally, we addressed the problem of evaluating the cost associated to producing the forecasts by calculating the Computing Time (CT). CT is defined as the time in seconds required for training and predicting the next $h$ time steps ahead with the model. Among the possible measures of complexity, such as the number of parameters, number of iterations, model’s depth, etc, CT is the most directly related to the monetary cost of forecasting since the forecasts are usually produced through some sort of cloud computing service that relies on pay-as-you-go fares (Spiliotis & Petropoulos, 2024). Hence, CT is analyzed for each forecasting model and retraining scenario combination. Lower CT values imply lower forecasting costs.
@@
Note that we related each evaluation metric result to the baseline retraining scenario, depending on the frequency of the dataset, to be able to easily compare the performance among different models and retrain windows. Furthermore, to statistically verify our research question, we tested the scenarios’ results using the Friedman-Nemenyi test for multiple comparisons (Demšar, 2006).